𝐴(1,βˆ’2) is a point on the circle (π‘₯βˆ’3)2+ (𝑦+1)2= 5

a. State the coordinates of the centre of the circle and hence find the coordinates of the point 𝐡 where 𝐴𝐡 is the diameter of the circle.

b. 𝐢(2,1) also lies on the circle. Use coordinate geometry to verify that angle 𝐴𝐢𝐡 = 900


1
Expert's answer
2022-02-07T17:49:19-0500

a.


"(\ud835\udc65\u22123)^2+ (\ud835\udc66+1)^2= 5"

The coordinates of the centre of the circle are "D(3, -1)."

If "AB" is the diameter of the circle, then "D" is the midpoint of "AB"


"x_D=\dfrac{x_A+x_B}{2}=>x_B=2x_D-x_A"

"y_D=\dfrac{y_A+y_B}{2}=>y_B=2y_D-y_A"

"x_B=2(3)-1=5, y_B=2(-1)-(-2)=0"

Point "B(5, 0)"


b.

Point "C(2, 1)"


"\overrightarrow{CA}=\langle1-2, -2-1\rangle"

"\overrightarrow{CB}=\langle5-2, 0-1\rangle"

"\overrightarrow{CA}\cdot \overrightarrow{CB}=-1(3)+(-3)(-1)=-3+3=0"

"|\overrightarrow{CA}|\not=0, |\overrightarrow{CB}|\not=0"

Therefore "\overrightarrow{CA}\perp \overrightarrow{CB}," and "\angle ACB=90\degree."